OpenStudy (anonymous):

A rectangular field is enclosed on 3 sides by a fence, while the 4th side is a stream. The fence opposite the stream is 12 m longer than the other sides. If the area of the field is 85 m2, find its dimensions.

OpenStudy (anonymous):

I have 3 more q's please someone help me with my last three so i can go to bed! thanks!

OpenStudy (anonymous):

Let short sides of fence = x Therefore third side of fence = x+12 Area of field = x(x+12) Therefore x(x+12)=85 x^2+12x=85 x^2+12x-85=0 (x+17)(x-5)=0 Therefore x=-17 or x=5 Since x>0, x=5 Therefore short side = 5m Third side of fence = 5+12=17m So dimensions of field are 5m by 17m